write the point-slope form of a line where m is the slope and (x1,y1) is a point on the line.
Answer by stanbon(48516)

write the point-slope form of a line where m is the slope and (x1,y1) is a point on the line.
------------
Point-slope form is y=mx+b
Need to find "b".
We know the slope is m and y=y1 when x=x1.
So you can solve for b as follows:
y1 = mx1 + b
b= y1-mx1
----------------
EQUATION:
y = mx + (y1-mx1)
y = m(x-x1) +y1
=================
Cheers,
Stan H.